OpenStudy (anonymous):

What is the equation of a line that passes through (7, 8) and has a slope of -3?

OpenStudy (anonymous):

y=mx+b M=slope x= x coordinate y= y coordinate Plug in your numbers, Solve for b Then replace the x and y coordinates with the variables x and y. :D

OpenStudy (anonymous):

Alternately, lines have an equation of form called "point-slope form"\[y-y_0 =m(x-x_0)\] which is handy because you can plug in your point (x_0 and y_0) and slope, then manipulate to whichever form is needed